Crew 219 Sol Summary Report 07-01-2020
Summary Title: Research, Research, Research
Author’s name: Dave Masaitis
Mission Status: Nominal
Sol Activity Summary: Sol 2 marks our full plunge into scientific research. With our mission plan published, we spent the day starting and continuing experiments, and conducted two more EVA’s for sample collection. Other features of the day include fresh bread and cobbler by the Lead Science Officer, and whispers of a showing of The Martian after the Comms window.
Look Ahead Plan: Tomorrow holds plans for two more proposed sample collection EVA’s, maintenance on a couple EVA suits, continued sample processing in the science dome, and continuation of our other experiments.
Anomalies in work: None
Weather: Cold and overcast all day, with temperatures below freezing.
Crew Physical Status: Nominal
EVA #3 – Geologic Sample Collection near North Ridge (0518600/4250900)
EVA #4 – Geologic Sample Collection around Kissing Camels (0518100/4249300)
Reports to be filed:
EVA Reports 1 & 2
EVA Requests 3 & 4
Photos (Including Photo of the Day)
(Carried over from previous day)
Sundries: Paper towels
Medical: C-Spine needs to be replaces (broken/misshapen)
12x Medium Gauze Pads